OpenStudy (anonymous):

4^x less than 64 answer in interval notation

OpenStudy (anonymous):

\[4^x < 64 \] 4^x = 64 x = log64/log4 x =3 x < 3

OpenStudy (anonymous):

\[4^x=64\].....\[4^x=4^3\]......\[x=3\]

OpenStudy (anonymous):

is it (0,3) or (-infinity,3)

OpenStudy (anonymous):

(-infinity,3)

OpenStudy (anonymous):

why